OpenStudy (anonymous):

Write aa polynomial function with rational coefficents so that P(x)=0 has the given roots. 1. 2,3,5 2. square root of 3, 2i

OpenStudy (ybarrap):

1 $$ P(x)=(x-2)(x-3)(x-5) $$ Because $$ P(2)=0\\ P(3)=0\\ P(5)=0 $$ Similarly, $$ P(x)=(x-\sqrt{3})(x-2i) $$ That's it!

OpenStudy (ybarrap):

Make sense?